if n(A)=30 ,n(B)=40 ,and n(AUB)=45,find n(A∩B)
Answer by Tatiana_Stebko(1539)

n(AUB)=n(A)+n(B)-n(A∩B)
45=30+40-n(A∩B)
n(A∩B)=70-45
n(A∩B)=25